OpenStudy (anonymous):

solve the equation x(5x+14)=3

OpenStudy (anonymous):

well its 5x ^2 + 14x = 3 ... use caltulator and solve :D whats hard on it??

OpenStudy (angela210793):

5x^2+14x-3=0 D=b2-4ac=196-4*5*(-3)=256 sqrD=16 2X1=(-b+-sqrD)/2a X1=1/5 and X2=3

OpenStudy (anonymous):

5x ^2 + 14x = 3 5x^2+14x-3=0 (5x-1)(1x+3)=0 x=1/5 or x=-3